# Local Order Average-Atom Interatomic Potentials

### v2

@@ III.1.3 Local Ordering Effect on Ground State RDF and Energy
To test the predictions of the phase diagram in Section III.1.2, we perform Monte Carlo (MC) simulations for the true species (TS) case (i.e. using the the species-dependent IP in Eq. (33)) to determine the ground state configuration for a 2D binary hexagonal crystal supercell. We begin with a crystal of 2688 all $A$ atoms with the equilibrium lattice constant of 1.1159. A specified concentration of $B$ atoms is then distributed onto the lattice. Simulations are performed for three different different choices of $\epsilon$ parameters corresponding to the two LO states and one no-LO state, and $\sigma=1$. To find the ground state configuration, MC swaps of atomic species are performed with a Metropolis acceptance criteria over a series of decreasing temperatures. 320,000 MC swaps are performed at each temperature, starting at $T=4.0$ (LJ temperature units) and halving each time until a final value of $T=1.953125\times 10^{-3}$ (see Supplementary Material (SM)). No relaxation is performed following MC swaps, i.e. the ideal all-$A$ crystal structure is held fixed. Since relaxation effects are small, this procedure leads to a good approximation for the true ground state. All simulations are performed using the LAMMPS package [lammps].
@@
The analysis in Section III.1.2 identifies three distinct regions in $\epsilon$ space that control the resulting ground state crystal structure as seen in Fig. 5 (right). $\epsilon$ values from the PS region of the phase diagram favor nearest-neighbor $B$ bonds, and as may be expected this leads to a ground state where the $B$ atoms clump together. Along the black line, no LO patterns are preferred, and the ground state shows a random pattern. Finally, in the SS region bonding between $B$ is unfavored, and in the ground state the $B$ atoms are spread apart.
Beyond visually validating the LO states, the results from the TS simulations are also examined in terms of their radial distribution functions. Partial RDFs for $AA$, ABAB, and $BB$ interactions from each region in the phase diagram are computed for first-, second-, and third-neighbor distances and shown in Fig. 6. The colors correspond to those in the phase diagram (see figure caption). The blue bars correspond to the total RDF of the system. The partial RDF results are averaged over 10 realizations with error bars shown. As expected from the derivation in Appendix A, the black bars of the randomly distributed system are equal to the total RDF. It can be seen that the three LO states exhibit distinct RDF profiles in accordance with their LO states. Interestingly, the partial RDFs for all points within the same region of the phase diagram are found to be equal, therefore Fig. 6 represents all partial RDFs for the ground state configurations of this system. This observation is explained by the no-LO line relation in Eq. (40). To the left of the no-LO line, $\epsilon_{BB}/\epsilon_{AA}>2\epsilon_{AB}/\epsilon_{AA}-1$, which means that there is a preference for $BB$ bonds since these are lower in energy and hence the same clumping of $B$ atoms is observed at all points in the PS region. In the SS region, the effect is opposite with $AB$ bonds preferred over $BB$ and hence the spreading of $B$ atoms.
Figure 6: Partial radial distribution functions at first, second, and third neighbors for (a) AAAA, (b) ABAB (c) BBBB interactions. Calculations performed for system size of $48\times 28$ unit cells ($N=2688$). Bar chart colors correspond to $\epsilon$ values from different regions of the phase diagram: green=PS, orange=$S$, and black=no-LO. The blue bars correspond to the total (non-partial) RDF. Partial RDFs shown are for $(\epsilon_{AB}/\epsilon_{AA},\epsilon_{BB}/\epsilon_{AA})=(1,3),(3,2),(1,1).$
We can further characterize the ordering seen in each LO state through the Warren-Cowley (WC) order parameters [cowley:1965, norman:warren:1951], defined by
$\alpha_{ij}^{XY}=1-\frac{p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C})}{c_{X}c_{Y}}$ (43)
for atoms $i$ and $j$ of species $X$ and $Y$, where the probability is defined in Eq. (14). [^13] The standard Warren-Cowley parameter formula involves a conditional probability, but it can trivially be shown that the given formulation with a joint probability is also correct [rao:curtin:2022]. Fig. 7 displays the WC parameters for each atom type combination at nearest neighbor distance. The three LO states are plotted in accordance with their coloring in Fig. 5. A WC parameter that is <0<0 indicates an attractive interaction between atom types, while >0>0 indicates repulsion. A WC parameter equal to 0 suggests that there is no LO, which explains why the no-LO state is not visible in Fig. 7. Only first neighbor WC parameters were plotted as the second neighbor parameters are nearly identical.
Figure 7: Warren-Cowley order parameters for nearest neighbor interactions. Calculations performed for system size of $48\times 28$ unit cells ($N=2688$). Colors correspond to those from phase diagram regions. $(\epsilon_{AB}/\epsilon_{AA},\epsilon_{BB}/\epsilon_{AA})=(1,3),(3,2),(1,1).$
Figure 8: TS energy contours over $\epsilon$-space. The no-LO line is shown in red.
The potential energy per atom across $\epsilon$ space is visualized in the energy contour plot in Fig. 8. The energy is computed at 100 points across phase diagram space. At each point, results are averaged over 10 ground state realizations obtained via MC with different random seeds. This plot can be understood by considering the different LO states on the phase diagram. In the SS region of Fig. 5, $B$ atoms are not (or are minimally) interacting, which explains the vertical energy contour lines in this region since there is little dependence on $\epsilon_{BB}$. In the PS region of the phase diagram $B$ atoms form clumps, and here we see that the energy has little dependence on $\epsilon_{AB}$. This makes sense since there are fewer $AB$ interactions as these are only present along the borders of the clumps of $B$ atoms.
@@ III.1.4 AA and LOAA Predictions for Ground State Energy
As seen from the TS analysis, multiple realizations of large and computationally expensive MC simulations are necessary to obtain reliable statistical estimates for the ground state properties. As explained in Section II, the AA and LOAA methods directly predict statistical averages with smaller supercells and without the need for repeated simulations. The AA approach assumes a random alloy neglecting the effect of LO, and LOAA includes LO effects via the partial RDFs of the ground state.
The AA potential form is given in Eqns. (8) and (9). For the the special case of a LJ pair potential where the $\sigma$ parameter for all interactions are the same, the effective AA potential energy has the simple form
@@
The LOAA potential is defined in Eqns. (16) and (17). The partial RDFs used in Eq. (17) are averaged over 10 different random seeds at 100 individual points across $\epsilon$ space using the TS simulations described in Section III.1.3. To implement the LOAA formulation, we create a new pair stlye in LAMMPS. The LOAA potential is created by modifying the existing LJ pair style to include the additional $G_{XY}(R)$ term seen in Eq. (16).
Figure 9: Difference in the energy per atom between the (a) AA method, and (b) LOAA method, and the exact TS results plotted across $\epsilon$-space. The no-LO line is shown in red.
The performance of the AA and LOAA is evaluated by comparing their predictions for the potential energy per atom with the exact results obtained using TS simulations across $\epsilon$-space.
The results for the AA method are presented in Fig. 9(a). As expected, we see that the method is exact only in the case of no LO (shown as a red line in the plot). The error in the AA energy increases linearly with the distance from this line as shown by a theoretical analysis in Appendix C. Comparing with the TS energy contour plot in Fig. 8, we find that the relative error in the AA energy prediction can be as large as 45% depending on the location in $\epsilon$-space (see Fig. S2 in the SM). Since, as explained in Section III.1.3, the partial RDFs are the same at all points in a given region of the phase diagram, the increasing error in AA method predictions is due to an increase in the $BB$ bond strength rather than an increase in the degree of LO.
We note that errors are larger to the left of the no-LO line (corresponding to the PS region in Fig. 5) relative to the right (SS region). Referring to the energy difference expression in Eq. (76), we see that the error in the direction $t$ perpendicular to the no-LO line scales with with $\zeta-2\gamma$. Due to the nature of the partial RDF profiles in the PS and SS regions, this factor is considerably larger in magnitude (around four times) in the PS region than in the SS region. A systematic increase in error occurs on both side of the no-LO line, but this explains why the error grows more rapidly with distance to the left of the no-LO line.
The results for the LOAA method are presented in Fig. 9(b) using the same contour scale as for the AA method in Fig. 9(a). The maximum error in energy for LOAA is about $6\times 10^{-3}$, not discernible on this scale. This small error reflects finite size limitations of the TS simulation rather than the LOAA formalism, which is exact. Both the partial RDFs and the energy per atom computed from the TS simulations were obtained from 10 realizations of a 2688 atom system, and as seen in Figs. 2 and 3, neither are fully converged. As a result both the TS energy and the LOAA-predicted energy (which depends on the partial RDFs) have errors and do not match exactly.
@@ III.1.5 AA and LOAA Predictions for Elastic Constants
To further validate the LOAA method, elastic constants are computed at various points in $\epsilon$-space using the TS, AA, and LOAA approaches. Similar to the ground state energy studies, we expect the AA method to be correct only for points on the no-LO line. In contrast, the LOAA method should yield accurate elastic constants for all $\epsilon$ values as it accounts for LO.
@@
Given the observation in Section III.1.4 that the error in AA prediction scales with the perpendicular distance $t$ from the no-LO line, we perform a similar analysis for the elastic constants. The results using each of the three methods (AA, LOAA, TS) are plotted in Figs. 10(a) and 10(b). The plots are nearly identical at all points along the no-LO line (i.e. at all values of $s$). We see that LOAA predictions agree precisely with the TS ground truth, whereas the AA predictions diverge in proportion to the distance from the no-LO line. The effect is particularly noticeable to the right of the no-LO line ($t>0$) corresponding to the SS region where there is a change in slope that AA fails to capture. This change in behavior is explained by an analytical derivation of the AA and LOAA elastic constants expressions in Appendix D. This analysis provides the scaling factors of the AA and LOAA elastic constants with distance from the no-LO line (i.e. the slopes in Figs. 10(a) and 10(b)). For example for $C_{11}$ the slopes are
We see that the slope for the AA method depends only on the concentrations $c_{A}$ and $c_{B}$ and the crystal structure and so will be the same in Regions I and II as seen in Fig. 10(a). In contrast the LOAA slope depends on the partial RDFs through the probability functions $\Phi_{AB}(r)$ and $\Phi_{BB}(r)$ (see Eq. (15)), which explains the change in slope across the no-LO line.
@@
As further tests and demonstrations of the LOAA method, we also explore two 3D EAM alloys. First, we consider the ordering effects on lattice constant, cohesive energy, and elastic properties in an FeNiCr alloy as described by the potentials of Bonny et al. [Bonny_2011]. This is the same system studied by [varvenne:luque:2016] for random alloys. We then consider temperature and stress dependent martensitic phase transformations in NiAl using the potential of [pun:mishin:2009].
All 3D simulations are performed with LAMMPS and visualized using Ovito.
@@ III.2.1 Ordering Effects in FeNiCr
In order to directly compare to the AA results of [varvenne:luque:2016], we consider alloys of the composition
Fe(1-x)/2Ni(1-x)/2Crx for various Cr concentrations $x$. As in the previous sections, we will denote the results of the full EAM description with atoms assigned their true species as “TS”. These can either be “random”, in which case atoms are randomly assigned a species consistent with the concentrations, or “ordered”, in which case the species of each lattice site is determined by MC as described below. All TS results are shown for a system of 32,00032{,}000 atoms on $20\times 20\times 20$ units cells of an fcc crystal with periodic boundary conditions.
@@
Results of the AA and LOAA formulations are compared to the TS results in Fig. 11. Note that both the AA and LOAA simulations can be performed on minimal periodic simulation cells (these were done on 32-atom fcc crystals). Since the formulations homogenize the atoms to make them all equivalent, there is no longer a need to represent the complex ordered (or disordered) state or to have enough atoms to accurately match the concentrations of each species. The additional requirement of the LOAA formulation — the function $G(R)$ — is determined from the RDFs of the final configuration of the same on-lattice MC anneal described above.
@@
III.2.2 Radial Distribution Functions for NiAl
Figure 13: (a)-(c) Partial RDFs for the ordered TS species, compared with the expected RDF for a random distribution of the Ni and Al atoms. (d) resulting $G(R_{0})$ for the Ni0.67Al0.33 alloy.
For the NiAl system, we begin by examining the partial RDFs, the resulting $G(R)$, and the effect these have on the effective pair potential, then in Section III.2.3 we examine NiAl phase transformations.
In Fig. 13, we compare the partial RDFs for the Ni–Ni, Ni–Al and Al–Al pairs in the ordered TS sample after the MC anneal (recall that this configuration is used to generate $G(R)$ for the LOAA effective pair potentials). In each case, we compare to the bcc RDF for the case where the atomic species are randomly assigned, which is just the RDF for ideal bcc ignoring species. These plots emphasize the effect of the ordering. For example, we see virtually no Al–Al bonds at the nearest neighbor distance, compared to the number of Ni–Ni and Ni–Al bonds. The LOAA formulation takes this into account through the $G(R)$ functions, as shown in Fig. 13(d). When considering a bond that is at some distance $R_{0}$ in the reference configuration, its contribution is a weighted sum of the three types of interactions, not only by average overall concentration (as in AA), but also by the relative probability of finding those bonding pairs in the ordered structure.
Figure 14: (a) Comparison of the TS pair potentials with the effective pair potential for the AA formulation, for the Ni0.67Al0.33 case. (b) Further comparison to the LOAA effective pair potentials, one for each of the first 5 non-zero values in the RDF functions.
The effective pair potentials for AA and LOAA are presented in Fig. 14. In AA, $V_{\rm eff}(r)$ is simply a blend of the three inter-species pair potentials, weighted by their respective compositions. In LOAA, each peak in the RDF curves, corresponding to a different reference configuration pair distance, $R_{0}$, produces a different effective pair potential, through the effect of $G(R)$. This is seen in Fig. 14(b), where the LOAA potential for the first five neighbor shell distances in the reference configuration are compared to the original EAM potentials and the blended result from AA. This highlights the pronounced differences between the AA and LOAA formulation, allowing LOAA to predict the stabilization of the B2 phase at high temperature.
III.2.3 Phase Transformations in NiAl
@@
Perhaps a more interesting and rigorous test of LOAA is the case of martensitic phase transformations in NiAl. Using the EAM potentials they developed [pun:mishin:2009], Pun and Mishin showed [pun:mishin:2010] that the B2 lattice is stable at high temperatures for a variety of off-stochiometric NiAl compositions, provided that the atomic species are determined through an appropriate MC annealing process (i.e. they are not simply random). Further, many of these compositions undergo a martensitic phase transformation to a face-centered tetragonal (fct) structure upon cooling for certain applied stresses. In the following we focus on the specific case of Ni0.67Al0.33   and consider, somewhat arbitrarily simply as an example, its phase transformations upon cooling from 900K to 100K and subsequent re-heating from 100K to 900K while subjecting it to a compressive stress of 900MPa along the $[001]$ lattice direction of the B2 structure.
@@
Such phase transformations will be, to some extent, stochastic in nature, and are therefore sensitive to the initial structure, initial velocities, system size, heating/cooling rate, and temperature control algorithm. Here, we make convenient choices and keep them consistent across all samples (with some clear exceptions explained below). Thus, we are comparing models (TS vs AA and LOAA) as opposed to trying to accurately predict transformation temperatures. Unless otherwise specified, we use a block of 6000 atoms in the B2 phase with an initial lattice constant of 2.8 Å. The box is initially orthogonal with the $x$, $y$ and $z$ axes oriented along $\bar{1}0] [1¯1$0][$\bar${1}10] and $[001]$, respectively. This is run using a Nosé–Hoover controlled NPT ensemble equilibrated at 900 K and with all stress components set to zero except the applied stress for 10,000 timesteps ($\Delta t=1$ fs). After this, the temperature is linearly ramped down from 900K to 100K over 500,000 steps, and then ramped back up to 900 over another 500,000 steps. This simulation allows the simulation box to change size and shape in order to maintain the correct stresses (it is a triclinic box in the language of LAMMPS, defined by three non-orthogonal periodic vectors). In the images of the simulation cell that follow, we show the $[001]$ or $z$ direction as “up” unless the image is specified as a TOP view, in which case the view is directly along the $z$-direction.
@@
By monitoring the length of the simulation box along the direction of the applied stress versus the temperature, we can capture the hysteresis loop of the phase transformations. Upon cooling, the length will contract slowly at first, due to simple thermal effects. It will then contract suddenly at a critical temperature called the martensite start temperature, $M_{s}$. Upon heating, there is again a sudden change ending on the original cooling line at the austenite finish temperature, $A_{f}$. Because $A_{f}$ is higher than $M_{s}$, a hysteresis loop is seen. This can be seen by skipping ahead to Fig. 16.
@@
Figure 15: (a) Initial structure of the random TS Ni0.67Al0.33 sample, colored by species (red is Ni, blue is Al). (b) The same structure as (a) colored by Ovito’s CNA, with blue indicating the B2 structure. (c) The final structure after a spontaneous phase transformation because the B2 structure is unstable. Here green is fcc-like and red is hcp-like, showing twin planes in the fcc-like fct crystal. (d) TOP view of (c) to show the degree of non-orthogonality after the transformation.
@@
First, we briefly consider the TS case where Ni and Al species are randomly assigned to a B2 lattice, as shown in Figs. 15(a) and (b). In (b), the atoms are coloured according to OVITO’s Common Neighbor Analysis (CNA) tool, which detects nominally bcc, fcc, and hcp local atomic arrangements. The B2 stucture of Ni0.67Al0.33 appears as nominally bcc (blue) using this tool. In this case, the B2 structure is not stable, and the crystal spontaneously transforms to a twinned fcc-like structure, seen in green and red in Fig. 15(c) and (d), where the twin planes appear as locally hcp-like and are colored red. The simulation cell remains nearly orthogonal during the transformation, but contracts along $z$ and expands along $x$ and $y$. This illustrates the importance of the LO (short-range order) in stabilizing the B2 structure. The random placement of the Al within the Ni lattice simply produces a random solid solution for which the bcc arrangement is not stable at any temperature, and the stable structure is fcc, distorted by the disparate sizes of the Ni and Al atoms.
@@
Figure 17: (a) Initial structure of the ordered TS Ni0.67Al0.33 sample, colored by species (red is Ni, blue is Al). (b) The same structure as (a) colored by Ovito’s CNA, with blue indicating the B2 structure. (c) The final structure after the cooling phase, which has transformed to a twinned fct structure. (c) The final structure upon reheating, restored to the original B2 with a few thermal fluctations.
Next, the TS case is run again, but this time with the atomic species set by an MC annealing process identical to the one described above for the FeNiCr system. The arrangement of the atomic species stabilizes the B2 structure at high temperature. The resulting hysteresis loop and structures are shown in Figs. 16 and 17. We take these plots to be the “ground truth” for the Ni0.67Al0.33 system — it is the curve we hope to reproduce with our averaged-atom models. In Fig. 17, we see that the crystal indeed transforms from the initial B2 in (a) and (b), to fct upon cooling in (c), and back to B2 after re-heating in (d). The fact that the low temperature phase is fct, and not fcc, is not obvious from the images, but can be verified by analyzing the lattice vectors carefully.
We now run the same simulation, but with the average-atom interatomic model (both AA and LOAA). We start with the same 6000 atom simulation box described above, but now all atoms are equivalent averaged-atoms instead of Ni or Al. We set their masses to be the geometric average based on the concentrations, i.e. $M=m_{Ni}^{0.67}m_{Al}^{0.33}=45.42$ amu, as per Section II.4.
Figure 18: (a) Final structure of the AA simulation after a spontaneous phase transformation because the B2 structure is unstable (compare to Fig. 15(c). (b) Alternate TOP view to highlight non-orthogonal box. (The initial structure was the same as Fig. 15(b).)

### v4

@@ III.1.3 Monte Carlo Simulations of the Ground State
To test the predictions of the phase diagram in Section III.1.2, we perform Monte Carlo (MC) simulations for the TS case (i.e. using the the species-dependent IP in Eq. (33)) to determine the ground state configuration for a 2D binary hexagonal crystal supercell. The simulations were performed for a crystal with concentrations $c_{A}=0.8$ and $c_{B}=0.2$. A convergence study (see Section I A in the Supplementary Material (SM)) indicates that a supercell containing 2688 atoms is sufficient. We begin with a crystal of all $A$ atoms. The specified concentration of $B$ atoms is then distributed onto the lattice. Simulations are performed for three different different choices of $\epsilon$ parameters corresponding to the two LO states and one no-LO state, and $\sigma=1$. MC swaps of atomic species are performed with a Metropolis acceptance criteria over a series of decreasing temperatures. 320,000 MC swaps are performed at each temperature, starting at $T=4.0$ (LJ temperature units) and halving each time until a final value of $T=1.953125\times 10^{-3}$ (see Section I B in the SM). No relaxation is performed following MC swaps, i.e. the ideal all-$A$ crystal structure is held fixed. Since relaxation effects are small, this procedure leads to a good approximation for the true ground state. All simulations are performed using the LAMMPS package [lammps].
@@
The analysis in Section III.1.2 identifies three distinct regions in $\epsilon$ space that control the resulting ground state crystal structure as seen in Fig. 2 (right). $\epsilon$ values from the PS region of the phase diagram favor nearest-neighbor $B$ bonds, and as may be expected this leads to a ground state where the $B$ atoms clump together. Along the black line, no LO patterns are preferred, and the ground state shows a random pattern. Finally, in the SS region bonding between $B$ is unfavored, and in the ground state the $B$ atoms are spread apart in a solid solution. (Partial RDFs and Warren-Cowley parameters to further validate the distinct LO states are included in Section I C in the SM.)
@@ III.1.4 AA and LOAA Predictions for Ground State Energy
As seen from the convergence analysis (Section I A in the SM), multiple realizations of large and computationally expensive MC simulations are necessary to obtain reliable statistical estimates for ground state properties like the cohesive energy (average energy per atom). In contrast, the AA and LOAA methods directly predict statistical averages with smaller supercells and without the need for repeated simulations. [^12] For LOAA a calculation of the partial RDFs is necessary, so no benefit is obtained unless multiple properties are computed.
The AA potential form is given in Eqns. (8) and (9). For the the special case of an LJ pair potential where the $\sigma$ parameter for all interactions are the same, the effective AA potential energy has the simple form
@@
The LOAA potential is defined in Eqns. (16) and (17). The partial RDFs used in Eq. (17) are averaged over 10 different random seeds at 100 individual points across $\epsilon$ space using the TS simulations described in Section III.1.3. To implement the LOAA formulation, we create a new pair style in LAMMPS. The LOAA potential is created by modifying the existing LJ pair style to include the additional $G_{XY}(R)$ term seen in Eq. (16).
Figure 3: TS cohesive energy contours over $\epsilon$-space. The no-LO line is shown in red.
The performance of the AA and LOAA methods is evaluated by comparing their predictions for the cohesive energy with the exact results obtained using TS simulations across $\epsilon$-space. The TS cohesive energy across $\epsilon$ space is visualized in the energy contour plot in Fig. 3. The energy is computed at 100 points across phase diagram space. At each point, results are averaged over 10 ground state realizations obtained via MC with different random seeds. (See Section I D in the SM for an explanation of the form of this plot.)
Figure 4: Difference in the cohesive energy between the (a) AA method, and (b) LOAA method, and the exact TS results plotted across $\epsilon$-space. The no-LO line is shown in red.
The difference between the AA and TS results are presented in Fig. 4(a). As expected, we see that the method is exact only in the case of no LO (shown as a red line in the plot). The error in the AA energy increases linearly with the distance from this line as shown by a theoretical analysis in Appendix D (see also the discussion in Section I E in the SM). Comparing with the TS energy contour plot in Fig. 3, we find that the relative error in the AA energy prediction can be as large as 45%.
The results for the LOAA method are presented in Fig. 4(b) using the same contour scale as for the AA method in Fig. 4(a). The maximum error in energy for LOAA is about $6\times 10^{-3}$, not discernible on this scale. This small error reflects finite size limitations of the TS simulation rather than the LOAA formalism, which is exact. Both the partial RDFs and the energy per atom computed from the TS simulations were obtained from 10 realizations of a 2688 atom system, neither of which are fully converged (see Figs. S2 and S3 in the SM). As a result both the TS energy and the LOAA-predicted energy (which depends on the partial RDFs) have errors and do not match exactly.
@@ III.1.5 AA and LOAA Predictions for Elastic Constants
To further validate the LOAA method, elastic constants are computed at various points in $\epsilon$-space using the TS, AA, and LOAA approaches. Similar to the ground state energy studies, we expect the AA method to be correct only for points on the no-LO line. In contrast, the LOAA method should yield accurate elastic constants for all $\epsilon$ values.
@@
Given the observation in Section III.1.4 that the error in AA prediction scales with the perpendicular distance $t$ from the no-LO line, we perform a similar analysis for the elastic constants. The results using each of the three methods (AA, LOAA, TS) are plotted in Figs. 5(a) and 5(b). The plots are nearly identical at all points along the no-LO line (i.e. at all values of $s$ (Fig. 15)). We see that LOAA predictions agree precisely with the TS ground truth, whereas the AA predictions diverge in proportion to the distance from the no-LO line. The effect is particularly noticeable to the right of the no-LO line ($t>0$) corresponding to the SS region where there is a change in slope that AA fails to capture. This change in behavior is explained by an analytical derivation of the AA and LOAA elastic constants expressions in Appendix E. The slope of the elastic constants for the AA method depends only on the concentrations $c_{A}$ and $c_{B}$ and the crystal structure and so will be the same in the PS and SS regions as seen in Fig. 5. In contrast the LOAA slope depends on the partial RDFs through the probability functions $\Phi_{AB}(r)$ and $\Phi_{BB}(r)$ (see Eq. (15)), which explains the change in slope across the no-LO line.
@@
As a further test of LOAA, we consider the more realistic case of 3D metallic alloys modeled via EAM potentials. First, we consider ordering effects on the lattice constant, cohesive energy, and elastic properties of Fe(1-x)/2Ni(1-x)/2Crx as described by the EAM potential of Bonny et al. [Bonny_2011]. This system was previously studied by Varvenne et al. [varvenne:luque:2016] treating it as a random alloy and applying the AA method. We then consider temperature and stress dependent martensitic phase transformations in Ni0.67Al0.33 using the EAM potential of Purja Pun and Mishin [pun:mishin:2009]. All 3D simulations are performed with LAMMPS and visualized using OVITO.
@@ III.2.1 Ordering Effects in FeNiCr
In order to directly compare to the AA results of [varvenne:luque:2016], we consider alloys of the composition Fe(1-x)/2Ni(1-x)/2Crx for various Cr concentrations $x$. As in the previous sections, we will denote the results of the full EAM description with atoms assigned their true species as “TS”. These can either be “random”, in which case atoms are randomly assigned a species consistent with the concentrations, or “ordered”, in which case the species of each lattice site is determined by MC as described below. All TS results are shown for a system of 32,00032{,}000 atoms on $20\times 20\times 20$ units cells of an fcc crystal with periodic boundary conditions.
@@
Results of the AA and LOAA formulations are compared to the TS results in Fig. 6. Note that both the AA and LOAA simulations can be performed on minimal periodic simulation cells (these were done on 32-atom fcc crystals). Since the formulations homogenize the atoms to make them all equivalent, there is no longer a need to represent the complex ordered (or disordered) state or to have enough atoms to accurately match the concentrations of each species. The additional requirement of the LOAA formulation — the functions $G_{XY}(R)$ (Eq. (17)) — are determined from the partial RDFs of the final configuration of the same on-lattice MC anneal described above.
@@
III.2.2 Phase Transformations in NiAl
@@
Next we consider the application of LOAA to martensitic phase transformations in NiAl. Using the EAM potentials they developed [pun:mishin:2009], Pun and Mishin showed [pun:mishin:2010] that the B2 lattice is stable at high temperatures for a variety of off-stoichiometric NiAl compositions, provided that the atomic species are determined through an appropriate MC annealing process (i.e. they are not simply random). Further, many of these compositions undergo a martensitic phase transformation to a face-centered tetragonal (fct) structure upon cooling for certain applied stresses. In the following we focus on the specific case of Ni0.67Al0.33   and consider, somewhat arbitrarily simply as an example, its phase transformations upon cooling from 900K to 100K and subsequent re-heating from 100K to 900K while subjecting it to a compressive stress of 900MPa along the $[001]$ lattice direction of the B2 structure.
@@
Such phase transformations will be, to some extent, stochastic in nature, and are therefore sensitive to the initial structure, initial velocities, system size, heating/cooling rate, and temperature control algorithm. Here, we make convenient choices and keep them consistent across all samples (with some clear exceptions explained below). Thus, we are comparing models (TS vs AA and LOAA) as opposed to trying to accurately predict transformation temperatures. Unless otherwise specified, we use a block of 6000 atoms in the B2 phase with an initial lattice constant of 2.8 Å. The box is initially orthogonal with the $x$, $y$ and $z$ axes oriented along $\bar{1}0] [1¯1$0][$\bar${1}10] and $[001]$, respectively. This is run using a Nosé–Hoover controlled NPT ensemble equilibrated at 900 K and with all stress components set to zero except the applied stress for 10,000 timesteps ($\Delta t=1$ fs). After this, the temperature is linearly ramped down from 900K to 100K over 500,000 steps, and then ramped back up to 900 over another 500,000 steps. This procedure allows the simulation box to change size and shape in order to maintain the correct stresses (it is a triclinic box in the language of LAMMPS, defined by three non-orthogonal periodic vectors). In the images of the simulation cell that follow, we show the $[001]$ or $z$ direction as “up” unless the image is specified as a TOP view, in which case the view is directly along the $z$-direction.
@@
By monitoring the length of the simulation box along the direction of the applied stress versus the temperature, we capture the hysteresis loop of the phase transformations. Upon cooling, the length will contract slowly at first, due to simple thermal effects. It will then contract suddenly at a critical temperature called the martensite start temperature, $M_{s}$. Upon heating, there is again a sudden change ending on the original cooling line at the austenite finish temperature, $A_{f}$. Because $A_{f}$ is higher than $M_{s}$, a hysteresis loop is seen. This can be seen by skipping ahead to Fig. 9.
@@
Figure 8: Snapshots of the random Ni0.67Al0.33 system computed using TS. (a) Initial structure colored by species (purple is Ni, cyan is Al). (b) The same structure as (a) colored by OVITO’s CNA, with blue indicating the B2 structure. (c) The final structure after a spontaneous phase transformation away from the B2 structure. Here green is fcc-like and red is hcp-like, showing twin planes in the fcc-like fct crystal. (d) TOP view of (c) to show the degree of non-orthogonality after the transformation.
@@
First, we briefly consider the TS case where Ni and Al species are randomly assigned to a B2 lattice, as shown in Fig. 8(a) and Fig. 8(b). In Fig. 8(b), the atoms are colored according to OVITO’s Common Neighbor Analysis (CNA) tool, which detects nominally bcc, fcc, and hcp local atomic arrangements. The B2 structure of Ni0.67Al0.33 appears as nominally bcc (blue) using this tool. In this case, the B2 structure is not stable, and the crystal spontaneously transforms to a twinned fcc-like structure, seen in green and red in Fig. 8(c) and Fig. 8(d), where the twin planes appear as locally hcp-like and are colored red. The simulation cell remains nearly orthogonal during the transformation, but contracts along $z$ and expands along $x$ and $y$. This illustrates the importance of the LO (short-range order) in stabilizing the B2 structure. The random placement of the Al within the Ni lattice simply produces a random solid solution for which the bcc arrangement is not stable at any temperature, and the stable structure is fcc, distorted by the disparate sizes of the Ni and Al atoms.
@@
Figure 10: Snapshots of the ordered Ni0.67Al0.33 system computed using TS. (a) Initial structure colored by species (purple is Ni, cyan is Al). (b) The same structure as (a) colored by OVITO’s CNA, with blue indicating the B2 structure. (c) The final structure after the cooling phase, which has transformed to a twinned fct structure. (c) The final structure upon reheating, restored to the original B2 with a few thermal fluctuations.
Next, the TS case is run again, but this time with the atomic species set by an MC annealing process identical to the one described above for the FeNiCr system. The arrangement of the atomic species stabilizes the B2 structure at high temperature. The resulting hysteresis loop and structures are shown in Figs. 9 and 10. We take these plots to be the “ground truth” for the Ni0.67Al0.33 system — it is the curve we hope to reproduce with the effective AA and LOAA models. In Fig. 10, we see that the crystal indeed transforms from the initial B2 in (a) and (b), to fct upon cooling in (c), and back to B2 after re-heating in (d). The fact that the low temperature phase is fct, and not fcc, is not obvious from the images, but can be verified by analyzing the lattice vectors carefully.
We now run the same simulation, but with effective IPs (both AA and LOAA). We start with the same 6000 atom simulation box described above, but now all atoms are equivalent averaged-atoms instead of Ni or Al. We set their masses to be the geometric average based on the concentrations, i.e. $M=m_{Ni}^{0.67}m_{Al}^{0.33}=45.42$ amu, as discussed in Section II.4.
Figure 11: Snapshots of the Ni0.67Al0.33 system computed using AA. (a) Final structure of the AA simulation after a spontaneous phase transformation because the B2 structure is unstable (compare to Fig. 8(c). (b) Alternate TOP view to highlight non-orthogonal box. (The initial structure was the same as Fig. 8(b).)